# uncompress

### v2

@@ 2.13. Relatively bounded families of couples
We define relatively bounded families of couples (respectively, varieties) as in [Bir23a, §3]. Let $\mathcal{P}$ be a family of couples $(X/Z,D)$. We say $\mathcal{P}$ is relatively bounded if there exist natural numbers $d,r$ such that for each $(X/Z,D)\in\mathcal{P}$, $\dim X-\dim Z\leq d$, there is a very ample/Z/Z divisor $A$ on $X$ such that
$\deg_{A/Z}A\leq r\penalty 10000\ \penalty 10000\ \mbox{and}\penalty 10000\ \penalty 10000\ \deg_{A/Z}D\leq r,$
and $f_{*}\mathcal{O}_{X}(A)$ is locally free where $f$ denotes $X\to Z$. The local freeness condition is to ensure that $A$ embeds $X$ into the projective bundle $\mathbb{P}(f_{*}\mathcal{O}_{X}(A))$. The local freeness is automatic when $Z$ is a smooth curve and this is the case needed in this paper.
When $D=0$ for every $(X/Z,D)\in\mathcal{P}$, we then refer to $\mathcal{P}$ as a family of relatively bounded varieties.
Here we collect some useful results about relatively bounded families of projective couples; for proofs, see [Bir23a, §3].
Lemma 2.14 (cf. [Bir23a, Lemma 3.2]).
@@
  • the horizontal/Z/Z part of $E$ is contained in $\operatorname{Supp}(G|_{Y})$ where $G|_{Y}$ is the divisorial part of the reduction of the closed subscheme $Y\times_{W}G$ of $Y$.
Then $\mathcal{P}$ is a relatively bounded set of couples.
@@
Now, for any given $(Y/Z,E)$, we can apply [Bir23a, Lemma 3.2] from some $1\leq i\leq r$. ∎
As for bounded set of projective couples (cf. [Bir19, Lemma 2.21]), there is a universal family of varieties and divisors for a relatively bounded family of projective couples over smooth curves.
Lemma 2.15 ([Bir23a, Lemma 3.4, Lemma 3.5]).
@@
    every component of $D$ is horizontal over $Z$, and
@@
Then there exist finitely many projective morphisms ${V}_{i}\to T_{i}$ of varieties and reduced divisors $C_{i}$ on $V_{i}$ such that for each $(X/Z,D)\in\mathcal{P}$ and each closed point $z\in Z$, after shrinking $Z$ around $z$ if necessary, there are an $i$ and a morphism $Z\to T_{i}$ such that $X=Z\times_{T_{i}}V_{i}$ and $D=Z\times_{T_{i}}C_{i}$.
@@ 2.16. Irrationality
Given a projective variety $X$, we define the degree of irrationality of $X$ following [BDELU17] as
$\operatorname{irr}(X)=\min\bigg\{\delta>0\,\,\bigg|\,\,\begin{array}[]{c}\exists\text{ degree }\delta\text{ rational dominant map}\\ X\dashrightarrow\mathbb{P}^{n}\text{ with }\dim X=n\end{array}\bigg\}.$
@@
By assumptions, we can assume that there is a single universal family $V\to T$ for $\mathcal{Q}$ as in §2.9. As irrationality is a birational invariant, we can assume that the varieties in $\mathcal{Q}$ are closed fibres of $V\to T$. Embed $V/T$ into a projective space $\mathbb{P}^{N}_{T}$ for some $N\in\mathbb{N}$. Take a general projection from $\mathbb{P}^{N}_{T}$ to some projective subspace of $\mathbb{P}^{N}_{T}$, then there exists a generically finite rational $T$-map $V\dashrightarrow\mathbb{P}^{n}_{T}$ for some $n\leq N$. Possibly taking a stratification of $\mathcal{Q}$ shows that the irrationality for $X\in\mathcal{Q}$ is bounded from above. ∎
@@ 3.1. Morphisms of couples
A morphism $(Z,E)\to(V,C)$ between couples is a morphism $f\colon Z\to V$ of schemes such that $f^{-1}(C)\subseteq E$.
@@ 3.2. Toric varieties and toric morphisms
An affine toric variety $X$ of dimension $d$ is an irreducible affine variety containing an algebraic torus $\mathbb{T}_{X}\cong\mathbb{G}_{m}^{d}$ as a Zariski open subset such that the action of $\mathbb{T}_{X}$ on itself extends to an algebraic action of $\mathbb{T}_{X}$ on $X$. When the affine toric variety $X$ is normal, this is equivalent to giving a pair $(N_{X},\sigma)$ where $N_{X}$ is a lattice of finite rank and $\sigma$ is a strongly convex rational polyhedral cone in $N_{X}\otimes_{\mathbb{Z}}\mathbb{R}$ such that $X=\operatorname{Spec}\mathbb{K}[\sigma^{\vee}\cap M_{X}]$ where $M_{X}$ is the dual lattice of $N_{X}$ and $\sigma^{\vee}\subset M_{X}\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$. A toric morphism of affine normal toric varieties $X$ and $Y$ is given by a linear map of lattices $\phi\colon N_{X}\to N_{Y}$ such that the image of the cone of $X$ under the extended map $\phi_{\mathbb{R}}\colon N_{X}\otimes_{\mathbb{Z}}\mathbb{R}\to N_{Y}\otimes_{\mathbb{Z}}\mathbb{R}$ is contained in the cone of $Y$. We refer to [CLS11] for the general theory of toric varieties.
If $D$ is the toric boundary of $X$, that is, $D$ is the reduction of complement of the big torus $\mathbb{T}_{X}$ of $X$, it is well-known that $(X,D)$ is lc and $K_{X}+D\sim 0$; cf. [CLS11, Theorem 8.2.3]. In this case, we say that $(X,D)$ is a toric couple.
@@ 3.3. Toroidal couples
of completions of local rings so that the ideal of $D$ is mapped to the ideal of the toric boundary divisor $C\subset W$, that is, the reduction of complement of the big torus $\mathbb{T}_{W}$ of $W$. We call $\{(W,C),w\}$ a local toric model of $\{(X,D),x\}$. We say the couple $(X,D)$ is toroidal if it is toroidal at every closed point.
In literature, the open immersion $U_{X}\coloneqq X\setminus\operatorname{Supp}D\subset X$ is called a toroidal embedding; for example, see [KKMS73, II §1]. We usually denote this toroidal embedding by $(U_{X}\subset X)$. Note that $U_{X}$ is smooth as $\mathcal{O}_{X,x}$ is regular if and only if $\widehat{\mathcal{O}}_{X,x}$ is regular. For citations, we will use the notions toroidal couples and toroidal embeddings interchangeably to be consistent with literature, for example, [AK00, KKMS73]. Moreover, if the embedding $(U_{X}\subset X)$, or equivalently the couple $(X,D)$, is clear from the context, we just say that $X$ is a toroidal variety.
Let $(X,D)$ be a toroidal couple. If every irreducible component of $D$ is normal, we call $(X,D)$ a strict toroidal couple. In this case, we say that the corresponding toroidal embedding $(U_{X}\subset X)$ is a strict toroidal embedding or a toroidal embedding without self-intersection; see [KKMS73, page 57].
Lemma 3.4 ([Bir23a, Lemma 3.8]).
@@
If $(X,D)$ is a toroidal couple, by Lemma 3.4, $K_{X}+D$ is Cartier, and $(X,D)$ is an lc pair, so $a(F,X,D)$ is a non-negative integer. Now, we assume that $(X,D)$ is a strict toroidal couple and that $a(F,X,D)=0$.
Step 1. Shrinking $X$ around the generic point of $\operatorname{centre}_{X}F$, by [KKMS73, page 195], there exist a normal affine toric variety $W$ and an étale morphism $p\colon X\to W$ such that $X\times_{Y}C$ is equal to $D$, where $C$ is the reduction of complement of the torus of $W$.
Let $X^{\prime}\to X$ be a log resolution such that $F$ is a nonsingular divisor on $X^{\prime}$. Denote by $\eta_{F}$ the generic point of $F$. By [Kol13, Lemma 2.22], there exists a morphism $W_{n}\to W$, which is the composite of a sequence of blowing ups, such that
@@
    there exists an induced rational map $f_{n}\colon X^{\prime}\dashrightarrow W_{n}$, and
@@
Since $p$ is étale, $X\times_{W}W_{n}$ is irreducible and smooth; moreover, the projection $X\times_{W}W_{n}\to X$ is birational. Then $f_{n}^{\prime}(\eta_{F})$ is also a codimension one regular point of $X\times_{W}W_{n}$, hence $f_{n}^{\prime}$ is an isomorphism near $\eta_{F}$ by [Liu06, Lemma 3.3.24]. Let $F_{n}$ be the closure of $f_{n}(\eta_{F})$. Then by [Kol13, (2.42.2)], $a(F_{n},W,C)=0$, that is, $F_{n}$ is a log canonical place of $(W,C)$.
Step 2. Let $\pi\colon W^{\prime}\to W$ be a toric log resolution of $W$, and let $\phi_{n}\colon W_{n}\dashrightarrow W^{\prime}$ be the induced birational map. If $C^{\prime}$ is the complement of the torus of $W^{\prime}$, then $K_{W^{\prime}}+C^{\prime}=\pi^{*}(K_{W}+C)$. Thus, the closure of $\phi_{n}(\eta_{F_{n}})$ in $W^{\prime}$, that is, $\operatorname{centre}_{W^{\prime}}(F_{n})$, is an lc centre of $(W^{\prime},C^{\prime})$, where $\eta_{F_{n}}$ is the generic point of $F_{n}$. Then $\operatorname{centre}_{W^{\prime}}(F_{n})$ is an irreducible component of the intersections of irreducible components of $C^{\prime}$ as $(W^{\prime},C^{\prime})$ is log smooth.
By the orbit-cone correspondence of toric varieties (see [CLS11, Theorem 3.2.6, Proposition 3.2.7]), each irreducible component of the intersections of prime divisors of $C^{\prime}$ is a disjoint union of finitely many orbits of the torus action, and every member of these irreducible components is a toric variety $V(\tau)=\overline{O(\tau)}$ for some orbit $O(\tau)$, where $\tau$ is a face of some cone in the fan defining $(W^{\prime},C^{\prime})$. Pick $V(\tau)$ so that $\operatorname{centre}_{W^{\prime}}(F_{n})=V(\tau)$. By blowing up $W^{\prime}$ along $V(\tau)$, we get a toric morphism $b\colon B_{V(\tau)}(W^{\prime})\to W^{\prime}$. Then $(B_{V(\tau)}(W^{\prime}),\lfloor b^{*}C^{\prime}\rfloor)$ is also log smooth and $\operatorname{centre}_{B_{V(\tau)}(W^{\prime})}(F_{n})$ is a log canonical centre of the pair $(B_{V(\tau)}(W^{\prime}),\lfloor b^{*}C^{\prime}\rfloor)$. We can blow up $\operatorname{centre}_{B_{V(\tau)}(W^{\prime})}(F_{n})$ further, and inductively the centre of $F_{n}$ on some blowing up $\widetilde{W}\to W^{\prime}$ will be a divisor generically isomorphic to $F_{n}$ by [Kol13, Lemma 2.22]. Since each step of the blowing ups is a toric morphism, we have $\widetilde{W}\to W^{\prime}$ is a toric morphism and $\widetilde{F}\coloneqq\operatorname{centre}_{\widetilde{W}}(F_{n})$ is a toric variety as a torus-invariant divisor on $\widetilde{W}$.
@@
Thus, the fibre $F^{\widetilde{\sigma}}_{\sigma}$ of $\pi_{\sigma}^{\widetilde{\sigma}}$ over a closed point in $O(\sigma)$ is isomorphic to the torus $\mathbb{T}_{\varphi^{-1}(N_{\sigma})/\widetilde{N}_{\widetilde{\sigma}}}$; see [HLY02, page 464]. Note that the fibre of $\pi^{\widetilde{\sigma}}_{\sigma}$ over a point of $O(\sigma)$ depends only on the orbit $O(\sigma)$; see [HLY02, Proposition 2.1.4].
By generic flatness, a general fibre of $\widetilde{F}\to V$ is pure dimensional of dimension $\dim\widetilde{F}-\dim V$, which is equal to $\dim O(\widetilde{\sigma})-\dim O(\sigma)$. Let $\widetilde{F}_{c}$ be a general fibre of $\widetilde{F}\to V$, where we can assume that $c\in O(\sigma)$. As $\dim\widetilde{F}_{c}\cap(\widetilde{F}\setminus O(\widetilde{\sigma}))<\dim\widetilde{F}_{c}$, there is an one-to-one correspondence between irreducible components of $\widetilde{F}_{c}$ and irreducible components of the open subset $O(\widetilde{\sigma})_{c}\cong F^{\widetilde{\sigma}}_{\sigma}\subset\widetilde{F}_{c}$. Therefore, a general fibre of $\widetilde{F}\to V$ is irreducible and rational.
Step 3. Since $F_{n}\dashrightarrow\widetilde{F}$ is a birational map of nonsingular varieties over $V$, a general fibre of $F_{n}\to V$ is also irreducible and rational. Denote by $F^{\prime}$ the closure of $f_{n}^{\prime}(\eta_{F})$, which is a prime divisor on $X\times_{W}W_{n}$; see Step 1. View $(\operatorname{centre}_{X}F)\times_{W}F_{n}$ as a closed subscheme of $X\times_{W}W_{n}$. As a general fibre of $F_{n}\to V$ is irreducible, $F^{\prime}$ is the only irreducible component of $(\operatorname{centre}_{X}F)\times_{W}F_{n}$ that dominates $\operatorname{centre}_{X}F$. Thus, a general fibre of $F^{\prime}\to\operatorname{centre}_{X}F$ is isomorphic to a general of $F_{n}\to V$. As $F\dashrightarrow F^{\prime}$ is a birational map over $\operatorname{centre}_{X}F$, by generic flatness, we can conclude that a general fibre of $F\to\operatorname{centre}_{X}F$ is irreducible and rational. ∎

### v3

@@ 2.13. Relatively bounded families of couples
We define (generically) relatively bounded families of couples (respectively, varieties) as in [11, §3]. Let $\mathcal{P}$ be a family of couples $(X/Z,D)$. We say $\mathcal{P}$ is generically relatively bounded if there exist natural numbers $d,r$ such that for each $(X/Z,D)\in\mathcal{P}$, we have the following: $\dim X-\dim Z\leq d$ and there is a very ample/Z/Z divisor $A$ on $X$ such that
$\deg_{A/Z}A\leq r~~\mbox{and}~~\deg_{A/Z}D\leq r.$
If in addition all the $(X/Z,D)\in\mathcal{P}$ are flat, we say that $\mathcal{P}$ is relatively bounded.
When $D=0$ for every $(X/Z,D)\in\mathcal{P}$, we then refer to $\mathcal{P}$ as a family of generically relatively bounded (respectively, relatively bounded) varieties.
Here we collect some useful results about (generically) relatively bounded families of couples; for proofs, see [11, §3].
Lemma 2.14 (cf. [11, Lemma 3.2]).
@@
  • the horizontal/Z/Z part of $E$ is contained in $\operatorname{Supp}(G|_{Y})$, where $G|_{Y}$ is the divisorial part of the reduction of the closed subscheme $Y\times_{W}G$ of $Y$.
Then $\mathcal{P}$ is a generically relatively bounded set of couples.
@@
Now for any given $(Y/Z,E)$, we can apply [11, Lemma 3.2] for some $1\leq i\leq r$. ∎
As for bounded set of projective couples (cf. [7, Lemma 2.21]), there is a universal family of varieties and divisors for a relatively bounded family of couples over smooth curves.
Lemma 2.15 ([11, Lemma 3.4, Lemma 3.5]).
@@
    every irreducible component of $D$ is horizontal over $Z$, and
@@
Then there exist finitely many projective morphisms ${V}_{i}\to T_{i}$ of varieties and reduced divisors $C_{i}$ on $V_{i}$ such that for each $(X/Z,D)\in\mathcal{P}$ and each closed point $z\in Z$, after shrinking $Z$ around $z$ if necessary, there exist an $i$ and a morphism $Z\to T_{i}$ such that $X=Z\times_{T_{i}}V_{i}$ and $D=Z\times_{T_{i}}C_{i}$.
@@ 2.16. Irrationality
Given a projective variety $X$, we define the degree of irrationality of $X$ following [6] as
$\operatorname{irr}(X)\coloneqq\min\bigg\{\delta>0\,\,\bigg|\,\,\begin{array}[]{c}\exists\text{ degree }\delta\text{ rational dominant map}\\ X\dashrightarrow\mathbb{P}^{n}\text{ with }\dim X=n\end{array}\bigg\}.$
@@
By assumptions, we can assume that there is a single universal family $V\to T$ for $\mathcal{Q}$ as in §2.9. As irrationality is a birational invariant, we can assume that the varieties in $\mathcal{Q}$ are closed fibres of $V\to T$. Embed $V/T$ into a projective space $\mathbb{P}^{N}_{T}$ for some $N\in\mathbb{N}$. Taking a general projection from $\mathbb{P}^{N}_{T}$ to some projective subspace of $\mathbb{P}^{N}_{T}$, there exists a generically finite rational map $\phi\colon V\dashrightarrow\mathbb{P}^{n}_{T}$ over $T$ for some $n\leq N$. There exists a nonempty open subset $U\subseteq T$ such that for each closed point $t\in U$, $\phi$ induces a generically finite rational map $\phi_{t}\colon V_{t}\dashrightarrow\mathbb{P}^{n}$ such that $\deg\phi_{t}=\deg\phi$, where $V_{t}$ is the fibre of $V\to T$ over $t$. The complement $T\setminus U$ is a union of varieties whose dimensions are strictly less than $\dim T$. Then proceeding by induction on $\dim T$, we can conclude that the irrationality of $X\in\mathcal{Q}$ is bounded from above. ∎
@@ 3.1. Morphisms of couples
A morphism $(Z,E)\to(V,C)$ between couples is a morphism $f\colon Z\to V$ of schemes such that $f^{-1}(C)\subseteq E$; see [11, §3.6].
@@ 3.2. Toric varieties and toric morphisms
An affine toric variety $X$ of dimension $d$ is an affine variety containing an algebraic torus $\mathbb{T}_{X}\cong\mathbb{G}_{m}^{d}$ as a Zariski open subset such that the action of $\mathbb{T}_{X}$ on itself extends to an algebraic action of $\mathbb{T}_{X}$ on $X$. When the affine toric variety $X$ is normal, this is equivalent to giving a pair $(N_{X},\sigma)$, where $N_{X}$ is a lattice of finite rank and $\sigma$ is a strongly convex rational polyhedral cone in $N_{X}\otimes_{\mathbb{Z}}\mathbb{R}$ such that $X=\operatorname{Spec}\mathbb{K}[\sigma^{\vee}\cap M_{X}]$, where $M_{X}$ is the dual lattice of $N_{X}$ and $\sigma^{\vee}\subset M_{X}\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$. A toric morphism of affine normal toric varieties $X$ and $Y$ is given by a linear map of lattices $\phi\colon N_{X}\to N_{Y}$ such that the image of the cone of $X$ under the extended map $\phi_{\mathbb{R}}\colon N_{X}\otimes_{\mathbb{Z}}\mathbb{R}\to N_{Y}\otimes_{\mathbb{Z}}\mathbb{R}$ is contained in the cone of $Y$. We refer to [16] for the general theory of toric varieties.
If $D$ is the toric boundary of $X$, that is, $D$ is the reduction of the complement of the torus $\mathbb{T}_{X}$ of $X$, it is well-known that $(X,D)$ is lc and $K_{X}+D\sim 0$; cf. [16, Theorem 8.2.3]. In this case, we say that $(X,D)$ is a toric couple.
@@ 3.3. Toroidal couples
of completions of local rings so that the ideal of $D$ is mapped to the ideal of the toric boundary divisor $C\subset W$, that is, the reduction of the complement of the torus $\mathbb{T}_{W}$ of $W$. We call $\{(W,C),w\}$ a local toric model of $\{(X,D),x\}$. We say the couple $(X,D)$ is toroidal if it is toroidal at every closed point, and we call $D$ the toroidal boundary.
In the literature, the open immersion $U_{X}\coloneqq X\setminus\operatorname{Supp}D\subset X$ is called a toroidal embedding; for example, see [32, II §1]. We usually denote this toroidal embedding by $(U_{X}\subset X)$. Note that $U_{X}$ is smooth as $\mathcal{O}_{X,x}$ is regular if and only if $\widehat{\mathcal{O}}_{X,x}$ is regular. For citations, we will use the notions of toroidal couples and toroidal embeddings interchangeably to be consistent with the literature, for example, [2, 11, 32]. Moreover, if the embedding $(U_{X}\subset X)$, or equivalently the couple $(X,D)$, is clear from the context, we just say that $X$ is a toroidal variety.
Let $(X,D)$ be a toroidal couple. If every irreducible component of $D$ is normal, we call $(X,D)$ a strict toroidal couple. In this case, we say that the corresponding toroidal embedding $(U_{X}\subset X)$ is a strict toroidal embedding or a toroidal embedding without self-intersection; see [32, page 57]. Moreover, if the toroidal boundary $D$, or equivalently the embedding $(U_{X}\subset X)$, is clear from the context, we just say that $X$ is a strict toroidal variety.
Lemma 3.4 ([11, Lemma 3.11]).
@@
If $(X,D)$ is a toroidal couple, by Lemma 3.4, $K_{X}+D$ is Cartier, and $(X,D)$ is an lc pair, so $a(F,X,D)$ is a non-negative integer. Now we assume that $(X,D)$ is a strict toroidal couple and that $a(F,X,D)=0$.
Step 1. After shrinking $X$ around the generic point of $\operatorname{centre}_{X}F$, by [32, page 195], there exist a normal affine toric variety $W$ and an étale morphism $p\colon X\to W$ such that $X\times_{W}C$ is equal to $D$, where $C$ is the reduction of the complement of the torus of $W$.
Let $X^{\prime}\to X$ be a log resolution such that $F$ is a nonsingular divisor on $X^{\prime}$. Denote by $\eta_{F}$ the generic point of $F$. By [34, Lemma 2.22], there exists a morphism $W_{n}\to W$, which is the composite of a sequence of blow-ups, such that
@@
    there exists an induced rational map $f_{n}\colon X^{\prime}\dashrightarrow W_{n}$, whose domain of definition contains a neighbourhood of $\eta_{F}$, and
@@
Since $p$ is étale, $X\times_{W}W_{n}$ is irreducible and smooth; moreover, the projection $X\times_{W}W_{n}\to X$ is birational. Then $f_{n}^{\prime}(\eta_{F})$ is also a codimension one regular point of $X\times_{W}W_{n}$, hence $f_{n}^{\prime}$ is an isomorphism near $\eta_{F}$ by [38, Lemma 3.3.24]. Let $F_{n}$ be the closure of $f_{n}(\eta_{F})$. Then by [34, §2.42], $a(F_{n},W,C)=0$, that is, $F_{n}$ is an lc place of $(W,C)$.
Step 2. Let $\pi\colon W^{\prime}\to W$ be a toric log resolution of $W$, and let $\phi_{n}\colon W_{n}\dashrightarrow W^{\prime}$ be the induced birational map. If $C^{\prime}$ is the reduction of the complement of the torus of $W^{\prime}$, then $K_{W^{\prime}}+C^{\prime}=\pi^{*}(K_{W}+C)$. Thus, the closure of $\phi_{n}(\eta_{F_{n}})$ in $W^{\prime}$, that is, $\operatorname{centre}_{W^{\prime}}(F_{n})$, is an lc centre of $(W^{\prime},C^{\prime})$, where $\eta_{F_{n}}$ is the generic point of $F_{n}$. Then $\operatorname{centre}_{W^{\prime}}(F_{n})$ is an irreducible component of the intersections of irreducible components of $C^{\prime}$ as $(W^{\prime},C^{\prime})$ is log smooth.
By the orbit-cone correspondence of toric varieties (see [16, Theorem 3.2.6, Proposition 3.2.7]), each irreducible component of the intersections of prime divisors of $C^{\prime}$ is a disjoint union of finitely many orbits of the torus action, and each of these irreducible components is a toric variety $V(\tau)=\overline{O(\tau)}$ for some orbit $O(\tau)$, where $\tau$ is a face of some cone in the fan defining $(W^{\prime},C^{\prime})$. Pick $V(\tau)$ so that $\operatorname{centre}_{W^{\prime}}(F_{n})=V(\tau)$. By blowing up $W^{\prime}$ along $V(\tau)$, we get a toric morphism $b\colon B_{V(\tau)}(W^{\prime})\to W^{\prime}$. Then $(B_{V(\tau)}(W^{\prime}),\operatorname{Supp}b^{*}C^{\prime})$ is also log smooth and $\operatorname{centre}_{B_{V(\tau)}(W^{\prime})}(F_{n})$ is a log canonical centre of the pair $(B_{V(\tau)}(W^{\prime}),\operatorname{Supp}b^{*}C^{\prime})$. We can blow up $\operatorname{centre}_{B_{V(\tau)}(W^{\prime})}(F_{n})$ further, and inductively the centre of $F_{n}$ on some blow-up $\widetilde{W}\to W^{\prime}$ will be a divisor generically isomorphic to $F_{n}$ by [34, Lemma 2.22]. Since each step of the blow-ups is a toric morphism, we have that $\widetilde{W}\to W^{\prime}$ is a toric morphism and $\widetilde{F}\coloneqq\operatorname{centre}_{\widetilde{W}}(F_{n})$ is a toric variety as a torus-invariant divisor on $\widetilde{W}$.
@@
Thus, the fibre $F^{\widetilde{\sigma}}_{\sigma}$ of $\pi_{\sigma}^{\widetilde{\sigma}}$ over a closed point in $O(\sigma)$ is isomorphic to the torus $\mathbb{T}_{\varphi^{-1}(N_{\sigma})/\widetilde{N}_{\widetilde{\sigma}}}$; see [28, page 464]. Note that the fibre of $\pi^{\widetilde{\sigma}}_{\sigma}$ over a closed point of $O(\sigma)$ depends only on the orbit $O(\sigma)$; see [28, Proposition 2.1.4].
By generic flatness, a general closed fibre of $\widetilde{F}\to V$ is pure-dimensional of dimension $\dim\widetilde{F}-\dim V$, which is equal to $\dim O(\widetilde{\sigma})-\dim O(\sigma)$. Let $\widetilde{F}_{c}$ be a general closed fibre of $\widetilde{F}\to V$, where we can assume that $c\in O(\sigma)$. As $\dim\big(\widetilde{F}_{c}\cap(\widetilde{F}\setminus O(\widetilde{\sigma}))\big)<\dim\widetilde{F}_{c}$, there is a one-to-one correspondence between irreducible components of $\widetilde{F}_{c}$ and irreducible components of the open subset $O(\widetilde{\sigma})_{c}\cong F^{\widetilde{\sigma}}_{\sigma}\subset\widetilde{F}_{c}$. Therefore, a general closed fibre of $\widetilde{F}\to V$ is irreducible and rational.
Step 3. Since $F_{n}\dashrightarrow\widetilde{F}$ is a birational map of nonsingular varieties over $V$, a general closed fibre of $F_{n}\to V$ is also irreducible and rational. Denote by $F^{\prime}$ the closure of $f_{n}^{\prime}(\eta_{F})$, which is a prime divisor on $X\times_{W}W_{n}$; see Step 1. View $(\operatorname{centre}_{X}F)\times_{W}F_{n}$ as a closed subscheme of $X\times_{W}W_{n}$. As a general closed fibre of $F_{n}\to V$ is irreducible, $F^{\prime}$ is the only irreducible component of $(\operatorname{centre}_{X}F)\times_{W}F_{n}$ that dominates $\operatorname{centre}_{X}F$. Thus, a general closed fibre of $F^{\prime}\to\operatorname{centre}_{X}F$ is isomorphic to a general closed fibre of $F_{n}\to V$. As $F\dashrightarrow F^{\prime}$ is a birational map over $\operatorname{centre}_{X}F$, by generic flatness, we can conclude that a general closed fibre of $F\to\operatorname{centre}_{X}F$ is irreducible and rational. ∎